OpenStudy (lilai3):

1) x to the negative 6th power over x to the negative 4th power 2) (y to the third power) to the negative 3rd power

OpenStudy (lgbasallote):

1) use \[\huge \frac{a^m}{a^n} = a^{m-n}\] 2) use \[\huge (a^m)^n = a^{mn}\]

OpenStudy (lilai3):

is the first one one over x to the second power?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

\[ 1) \quad \frac{1}{x^2} \checkmark\]

OpenStudy (lilai3):

kewl okay thanks is the second one one over y to the power of 6

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

remember \[\huge (a^m)^n \implies a^{mn} \quad \textbf{NOT} \quad a^{m+n}\]

OpenStudy (lilai3):

oh wait mistake...\[(^{y3})^{-2}\]

OpenStudy (lgbasallote):

oh..now you're right

OpenStudy (lgbasallote):

\[2) \quad \frac{1}{x^6} \checkmark\]

OpenStudy (lgbasallote):

that's a y

OpenStudy (lilai3):

yeah so number one and two are correct right?

OpenStudy (lgbasallote):

yup

OpenStudy (lilai3):

thank you

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (lilai3):

uhm i also have another question x to the 7th pwer times x to the 5th power

OpenStudy (lilai3):

(i'm checking my hmwk so yeah)